Circular No. 7622

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001bh IN UGC 11635 D. W. Xu and Y. L. Qiu, on behalf of Beijing Astronomical Observa-

tory supernova survey, report the discovery of an apparent supernova found on unfiltered CCD images taken on May 8.77 (poor image quality, magnitude not estimated) and 10.83 UT (mag ∼ 17.0). SN 2001bh is located at α = 20^{h}43^{m}29^{s}.82, δ = +80^{o}09′27′′.3 (equinox 2000.0), which is 1′′.8 west and 6′′.4 north of the nucleus of UGC 11635. A CCD image taken on 2000 Oct. 16.48 (limiting mag 18.2) showed nothing at this position.

SUPERNOVA 2001bg IN NGC 2608 A. Gal-Yam and O. Shemmer, Tel Aviv University, report: “A spec-

trum (range 400–780 nm) of SN 2001bg (cf. IAUC 7621) was obtained by J. Dann using the Wise Observatory 1-m telescope (+ FOSC spectrograph) on May 10. The spectrum is similar to early spectra of the type-Ia SN 1994D before maximum, including the prominent silicon feature redward of 600 nm.”

H. Kawakita and K. Kinugasa, Gunma Astronomical Observatory (GAO); and H. Yamaoka, Kyushu University, write: “A low-resolution spectrum (range 380–750 nm) of SN 2001bg was taken with GAO 0.65m telescope on May 10.55 UT. The preliminary reduction shows that the object to be a type-Ia supernova near maximum light. Adopting the NED recession velocity of 2135 km/s for NGC 2608, the SN expansion velocity, measured from the Si II absorption minimum (rest 635.5 nm), is 9700 km/s. Note that NGC 2608 also produced SN 1920A (m_{pg} = 11.8).”

SUPERNOVA 2001G IN MCG +08-17-43 T. Cameron reports the following CCD magnitudes of SN 2001G ob-

tained at the University of Calgary’s 1.8-m telescope at the Rothney Astrophysical Observatory: V filter, Jan. 23.18 UT, 14.65 ± 0.005; 30.14, 14.90 ± 0.007; Feb. 3.18, 15.09 ± 0.010; 17.15, 15.46 ± 0.011; 21.15, 15.76 ± 0.007; 27.14, 16.14 ± 0.012; Mar. 4.14, 16.28 ± 0.011; 20.18, 16.65 ± 0.018. B filter, Jan. 23.19, 14.49 ± 0.005; 30.15, 14.20 ± 0.005; Feb. 3.19, 15.10 ± 0.007; 17.18, 15.86 ± 0.010; 21.16, 16.47 ± 0.025; 27.15, 16.91± 0.017; Mar. 4.19, 16.95 ± 0.025. R filter, Jan. 23.20, 14.10 ± 0.006; 30.16, 14.34 ± 0.007; Feb. 3.20, 14.60 ± 0.009; 17.17, 14.85 ± 0.010; 27.16, 15.30 ± 0.011; Mar. 4.16, 15.57 ± 0.013.

2001 May 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green